### v1

marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style.Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
Native Parallel Reasoner: Reasoning in Parallelism via
Self-Distilled Reinforcement Learning
Tong Wu $^{\dagger}$ Yang Liu*  Jun Bai*  Zixia Jia Shuyi Zhang  Ziyong Lin  Yanting Wang
[…]
NLCo Lab, Beijing Institute for General Artificial Intelligence (BIGAI)
[…]
Figure 1: Native Parallel Reasoner (NPR) transforms a base model from sequential chain-of-thought (CoT) to native parallel reasoning via a self-distilled progressive training paradigm. Compared with previous SoTA, NPR achieves high reasoning accuracy, genuine parallelism and token acceleration. The illustrated results are evaluated on the AIME25 benchmark.
1 Introduction
The advent of super-scale Large Language Models (LLMs), exemplified by Gemini 3 (gemini3), GPT-5 (gpt5), and DeepSeek-V3.2 (deepseek_v3.2), has shifted the frontier of AI from semantic fluency to deep, multi-step agentic reasoning. Despite the excitement of “deeper” test-time scaling that enables models to solve complex problems muennighoff2025s1simpletesttimescaling, the “wider” reasoning capacity to explore diverse trajectories in parallel emerges as the dominant requirement toward agentic AI (shen2025thinkingvsdoingagents, gemini2.5). The MapReduce paradigm has long underpinned distributed computing by separating task decomposition from trajectory aggregation (mapreduce, multiverse, wang2025llmtimesmapreducev2), yet its application to agentic language modeling remains a critical missing link in the evolution of open-source LLMs. Ideally, the model should internalize the collaborative breadth of multi-agent systems directly into an efficient, natively parallel architecture.
Despite this clear imperative, existing implementations remain fragmented and present three critical deficiencies. First, Algorithmic and Architectural Incompatibility. Prevalent inference engines (sglang, vllm) and Reinforcement Learning (RL) algorithms (grpo) are ill-equipped for native branching: The former fails to control parallel branching and aggregation; the latter often clips the gradients of the special tokens that trigger those operations, preventing the model from learning strict structure. Second, Inefficient Hand-Crafted Parallelism. Although the intuitive advantage of parallel sampling lies in efficiency, early attempts on internalizing the parallelism parallel_r1, DBLP:journals/corr/abs-2509-04475/parathinker, zhao2025parallelsearch resort to hand-crafted divide-and-conquer rules via independent sampling. These methods fail to leverage shared Key-Value (KV) states, necessitating redundant recalculations for every branch and resulting in prohibitive linear latency costs ($O(N)$) that render the model impractical for real-time deployment. Third, Reliance on Supervised Distillation. Framework such as Multiverse (multiverse) successfully operationalizes native parallelism but depends heavily on supervised data distilled from stronger teacher models. While effective for compressing capabilities into smaller models, this dependence restricts the student to mimicking the teacher’s sequential reasoning topology force-fitted into a parallel format, imposing an "Intelligence Ceiling" that prevents it from novel, model-intrinsic parallel strategies necessary for super-intelligence.
To address these challenges, we take the first step to explore the potential of LLMs to self-evolve parallel reasoning capabilities without reliance on external supervision and introduce Native Parallel Reasoner (NPR). Specifically, NPR employs a three-stage progressive training paradigm designed to transition the model from sequential emulation to genuine parallel cognition. In Stage 1 ($\S$2.2), we warm up a seed instruction-tuned model (e.g., Qwen3-4B-Instruct) to spontaneously discover valid parallel structures by applying standard DAPO (dapo) with a format-aware reward function, yielding a structured trajectory generator, NPR-Zero, which produces parallel-formatted outputs but still relies on sequential visibility, i.e., simulated parallelism. In Stage 2 ($\S$2.3), we bridge the gap to native parallel architecture by performing rejection sampling on NPR-Zero and conducting parallel warmup, which instills strict topological constraints using parallel positional encoding and attention mask as in multiverse, converting the sequential behavior of models into real parallel execution. Finally, Stage 3 ($\S$2.4) generalizes these capabilities beyond the initial distribution via Native-Parallel RL. This stage implements collision-free parallel rollouts using a novel Parallel-Aware Policy Optimization (PAPO) algorithm, which optimizes branching policies directly within the parallel execution graph, allowing the model to learn adaptive decomposition strategies through trial and error rather than imitation.
To support this algorithmic breakthrough, we re-engineered the rollout infrastructure with a robust NPR Engine ($\S$2.5). Specifically, we observed several stability issues unique to parallel RL, e.g., GPU memory leaks caused by radix-cache mechanism; excessively long generation due to the incorrect parallel token calculations; potential runtime failure because of adaptive parallel inference logic, etc. NPR Engine redesigns memory management and flow control, providing the first stable rollout backend capable of supporting large-scale parallel RL.
We experiment NPR on Qwen3-4B-Instruct-2507 and Qwen3-4B (Non-Thinking) across a broad suite of eight reasoning benchmarks, demonstrating consistent performance improvements on both original model and previously RL-tuned version up to 24.5% and inference speedup up to 4.6×. The results reveal three consistent superiorities of NPR ($\S$3.3).
  • Self-Distilled Data Efficacy: Our self-distilled datasets outperform previous teacher-generated trajectories in multiverse by an average of 10.1 points, validating the hypothesis to learn from native distributions.
  • Parallelism Effectiveness and Efficiency: Both NPR-Beta and NPR-RL yields significant performance gains over direct sequential RL baselines (e.g., DAPO), confirming that adaptive parallel policies provide a superior search mechanism compared to single-path rollouts.
  • 100% Genuine Parallelism: We observed 30%+ AR fallback on test cases when running previous baselines ($\S$4.2), where models choose run vanilla AR generation to reach better performance. In contrast, NPR performs 100% genuinely parallel reasoning, with no instances of hidden AR fallbacks or pseudo-parallel behavior in all evaluated test cases.
Finally, we conduct comprehensive analyses of NPR spanning inference acceleration (§4.1), test-time scalability (§4.3), pseudo-parallelism (§4.2), evolution dynamics (§4.4), and qualitative case studies (§4.5). NPR achieves task-dependent speedups, reaching up to 4.6× over autoregressive (AR) decoding. Using best@8 as the metric for test-time scalability, we find that both parallel SFT and parallel RL consistently boost the performance of best-case exploitation across most benchmarks. Our qualitative studies highlight how NPR adapts its degree and style of parallelism across problem types, illustrating how structured parallel exploration leads to both faster inference and higher solution reliability.
1.1 Contributions
[…]
  • We introduce a teacher-free framework that unifies self-distilled data construction, parallel SFT, and parallel RL to build a truely NPR. Our method intrinsically learns adaptive decomposition, diverse parallel plans, and KV-reusing execution policies—establishing a reusable cognitive primitive.
  • We introduce Parallel-Aware Policy Optimization (PAPO), an RL algorithm specifically designed to optimize parallel decoding policies. Our analysis confirms that PAPO induces genuine parallel behaviors, where the model actively leverages independent attention branches for exploration and self-correction. Through case studies, we observe that NPR’s parallelism manifests not only through task decomposition but also through diverse exploratory branches and reflective cross-checks across different reasoning angles.
[…]
  • We re-engineered an NPR Engine, which redesigns core components of parallel execution, including radix-cache memory management, output-length control, and parallel-state branching based on Multiverse-Engine (multiverse). This eliminates the instability inherent in standard engines, providing a practical backend for native parallel RL training.
[…]
  • Using self-distilled parallel data, NPR reached even better performance gains. NPR-4B finetuned on Qwen3-4B-Instruct achieves 50.4% on AIME 25, 63.3% on AIME 24, surpassing the powerful baselines including Multiverse-4B(+14.9%) and Multiverse-32B(+12.5%) trained on distilled data s1.1. When training on Qwen3-4B (Non-thinking) [^1] We technically observe Qwen3-4B (Thinking) to be challenging finetuning its thinking tokens into desired formats. (§4.6) , NPR reaches a average performance gain of over 24.5%.
[…]
Figure 2: An overview of the NPR training framework.
2 Native Parallel Reasoner
In this study, we propose Native Parallel Reasoning (NPR), a framework that enables language models to generate and evaluate multiple reasoning branches in parallel. As shown in Figure˜2, NPR is developed through a three-stage curriculum that progressively induces, grounds, and amplifies this capability. First, NPR-Zero uses reinforcement learning to induce a structured parallel format without relying on external annotations. Next, NPR-Beta stabilizes these emerging parallel primitives through supervised fine-tuning on self-distilled trajectories. Finally, NPR applies a parallel-aware reinforcement learning procedure that directly optimizes the model’s ability to perform native parallel reasoning. Together, these stages establish a cohesive path from initial format induction to fully optimized parallel inference.
2.1 Preliminaries
Parallel Reasoning.
Parallel Reasoning (PR) relaxes the strict left-to-right dependency of AR reasoning, allowing the model to generate multiple reasoning steps independently whenever possible. Formally, the joint probability of a reasoning sample $\hat{y}$ consisting of $T$ reasoning steps $\{s_{t}\}_{t=1}^{T}$ can be factorized according to a dependency graph $\mathcal{G}$ defined over the steps:
$P(\hat{y}\mid q;\theta)=\prod_{t=1}^{T}P(s_{t}\mid\text{Pa}(s_{t}),q;\theta),$
where $\text{Pa}(s_{t})$ denotes the set of parent steps that $s_{t}$ directly depends on in $\mathcal{G}$, and $\theta$ are the model parameters. This formulation enables the model to process reasoning steps that do not have mutual dependencies concurrently.
Policy Optimization for Language Models
To optimize the policy model within our reinforcement learning framework, we adopt objective functions based on DAPO (dapo). We first introduce the original DAPO update procedure. For each question-answer pair $(q,y)\sim\mathcal{D}$, the policy model $\pi_{\theta_{\text{old}}}$ first generates a group of responses $\{\hat{y}_{i}\}_{i=1}^{G}$. The objective function $\mathcal{J}(\theta)$ is then formulated as:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{(q,y)\sim\mathcal{D},\,\{\hat{y}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}$ (1)
$\quad-\frac{1}{\sum_{i=1}^{G}|\hat{y}_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|\hat{y}_{i}|}\Bigg[\min\!\Bigg(r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\!\left(r_{i,t}(\theta),1-\epsilon_{\text{low}},1+\epsilon_{\text{high}}\right)\hat{A}_{i,t}\Bigg)\Bigg].$
$\quad\text{s.t.}\quad 0<\bigl|\{\hat{y}_{i}\mid\text{is\_equivalent}(y,\hat{y}_{i})\}\bigr|<G$
where $r_{i,t}(\theta)$ denotes the probability ratio between the current and the old policy for the $t$-th token in response $\hat{y}_{i}$, and $\hat{A}_{i,t}$ represents the standardized advantage of that token computed from the rewards $\{R_{1},R_{2},\ldots,R_{G}\}$ of all generated responses in the group:
$r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(\hat{y}_{i,t}\mid q,\hat{y}_{i,<t})}{\pi_{\theta_{\text{old}}}(\hat{y}_{i,t}\mid q,\hat{y}_{i,<t})},\quad\hat{A}_{i,t}=\frac{R_{i}-\text{mean}\!\left(\{R_{1},R_{2},\cdots,R_{G}\}\right)}{\text{std}\!\left(\{R_{1},R_{2},\cdots,R_{G}\}\right)}.$ (2)
This formulation ensures stable policy updates by clipping extreme probability ratios while encouraging exploration through group-wise normalization of advantages. It effectively balances between exploiting high-reward responses and maintaining diversity among generated outputs.
Table 1: Structured schema of NPR.
The Output Format Example of Parallel Reasoning
<guideline>
<plan>1: [One-sentence independent strategy]</plan>
<plan>2: [One-sentence independent strategy]</plan>
</guideline>
<step>1: [Self-contained detailed analysis for plan 1]</step>
<step>2: [Self-contained detailed analysis for plan 2]</step>
<takeaway>[Compare steps, synthesize findings, determine next action]</takeaway>
<guideline>
<plan>1: [One-sentence strategy]</plan>
</guideline>
<step>1: [Self-contained detailed analysis]</step>
<takeaway>[Final synthesis and conclusion]</takeaway>
[…]
2.2 Stage 1: Format-follow Reinforcement Learning
To support adaptive decomposition and parallel reasoning during generation, we adopt a simplified “Map–Process–Reduce” schema inspired by Multiverse (multiverse) but with a leaner structure. Each parallel block begins with <guideline> ... </guideline>, which contains a set of <plan> ... </plan> entries that define the Map stage. The Process stage follows: each <step> ... </step> block executes one mapped subtask independently and in parallel. After all <step> blocks complete, a Reduce stage consolidates their outputs into a final summary wrapped by <takeaway> ... </takeaway>. This explicit tag-based format makes the decomposition, independent processing, and final aggregation easy to parse and verify in downstream training and evaluation.
While this schema provides a clear, learnable format for parallel reasoning, obtaining large-scale, high-quality training data for it remains challenging. Prior work such as Multiverse (multiverse) constructs large, multi-step synthetic pipelines and aggregates outputs from several state-of-the-art teacher models (e.g., Deepseek R1 (deepseek_r1) and Gemini 2.5 Pro (gemini2.0)) to overcome data scarcity. While effective, these multi-teacher pipelines add operational complexity, require access to strong external teachers, and incur substantial maintenance costs.
We adopt a simpler, self-improving approach. Starting from a single pretrained LLM, we apply DAPO (dapo) to induce the target native parallel-reasoning generation format without paired supervision or external teachers. Our reward function combines format and accuracy signals. For format: outputs that pass a format check receive a reward of 0.0; outputs that fail receive a penalty in (0.0, -2.0]. For accuracy: when the format check passes, correct answers yield +1.0 and incorrect answers yield -1.0. The checkpoint produced by this process (denoted NPR-Zero) is therefore optimized primarily to learn the required structured format; we then use its generations for large-scale self-distillation to build a synthetic corpus for downstream supervised fine-tuning (SFT).
This pipeline removes the dependency on multiple external teacher models and produces a scalable, structured dataset that supports subsequent SFT stages.
2.3 Stage 2: Rejection Sampling and Parallel Warmup
Algorithm 1 Parallel Attention Mask
1:sequence: $\mathcal{I}:=\{t_{1},\dots,t_{L}\}$;
2:      Tag tokens: $\{\tau_{\text{parallel}}^{\pm},\tau_{\text{step}}^{\pm},\tau_{\text{plan}}^{\pm}\}$.
3:Attention mask: $\mathbf{M}\in\mathbb{R}^{L\times L}$.
4:procedure Construct NPR Attn Mask
5:  $\mathbf{M}\leftarrow\text{tril}(\mathbf{1}_{L\times L})$ $\triangleright$ Causal mask
6:  $\mathcal{S}\leftarrow\emptyset$ $\triangleright$ Init structure stack
7:  for $i=1\dots L$ do
8:   if $t_{i}\in\{\tau_{\text{parallel}}^{+},\tau_{\text{step}}^{+},\tau_{\text{plan}}^{+}\}$ then
9:     $\mathcal{S}.\text{push}(\{\text{type}(t_{i}),i\})$
10:   else if $t_{i}\in\{\tau_{\text{step}}^{-},\tau_{\text{plan}}^{-}\}$ then
11:     $b\leftarrow\mathcal{S}.\text{pop}()$
12:     Save span $(b.\text{start},i)$ in parent block
13:   else if $t_{i}=\tau_{\text{parallel}}^{-}$ then
14:     $b\leftarrow\mathcal{S}.\text{pop}()$
15:     $\{\mathcal{P}_{j}=[s_{j},e_{j})\}_{j=1}^{n}\leftarrow b.\text{steps}$
16:     for $(j,k)\in[1,n]^{2}$ where $j\neq k$ do
17:      $\mathcal{I}_{j}\leftarrow\{s_{j},\dots,e_{j}-1\}$
18:      $\mathcal{I}_{k}\leftarrow\{s_{k},\dots,e_{k}-1\}$
19:      $\mathbf{M}[\mathcal{I}_{j},\mathcal{I}_{k}]\leftarrow 0$ $\triangleright$ Isolate steps
20:      $\mathbf{M}[\mathcal{I}_{k},\mathcal{I}_{j}]\leftarrow 0$
21:  $\mathbf{M}\leftarrow\begin{cases}0&\text{if }\mathbf{M}[i,j]=1\\ -\infty&\text{if }\mathbf{M}[i,j]=0\end{cases}$
22:  return $\mathbf{M}$
Structured Trajectories Collection via Rejection Sampling.
To obtain high-quality structured reasoning traces without relying on external annotations, we employ a simple self-distillation procedure. For each question $q_{i}\in\{q_{1},q_{2},\dots,q_{N}\}$ in the dataset, the model generates $K$ candidate reasoning trajectories and corresponding answers $\{(r^{i}_{j},\hat{a}^{i}_{j})\}_{j=1}^{K}$ by repeated sampling. These samples form the pool from which we extract positive supervision signals.
We apply a rejection-sampling filter designed to mirror the bootstrapping setup used in NPR-Zero. Each sampled trajectory is evaluated using two lightweight, indicator-style constraints:
[…]
  • Structured Parallelism: To ensure clean supervision for parallel generation, we remove any trajectory that fails to adhere to the required structured output format (Table˜1). This constraint is encoded as $\mathbb{1}_{\text{format}}(r)$.
A sample is accepted only if it satisfies both criteria:
$\mathbb{1}_{\text{accept}}(r,\hat{a})=\mathbb{1}_{\text{correct}}(\hat{a})\cdot\mathbb{1}_{\text{format}}(r).$ (3)
Applying this filter yields the distilled dataset
$\begin{split}\mathcal{D}_{\text{accept}}=\{(q_{i},r^{i}_{j},a^{i}_{j})\mid i\leq N,j\leq K,\text{ s.t. }(r^{i}_{j},a^{i}_{j})\sim\pi_{\theta}(\cdot|q_{i}),\mathbb{1}_{\text{accept}}(r^{i}_{j},\hat{a}^{i}_{j})=1\}.\end{split}$ (4)
These accepted trajectories serve as the training corpus for the subsequent supervised fine-tuning stage, which provides a stable initialization for the parallel RL procedure described in §2.4.
Algorithm 2 Parallel Positional Encoding
1:Token sequence $\mathcal{I}:=\{t_{1},\dots,t_{L}\}$; Tag tokens $\{\tau_{\text{parallel}}^{\pm},\tau_{\text{step}}^{\pm},\tau_{\text{guideline}}^{\pm}\}$
2:Position IDs: $\mathbf{P}\in\mathbb{R}^{L}$
3:procedure Construct NPR Position IDs
4:  $\mathbf{P}\leftarrow[0,1,\dots,L-1]$; $\mathcal{S}\leftarrow\emptyset$ $\triangleright$ Init sequential positions & block stack
5:  for $i=1\dots L$ do
6:   $b\leftarrow\mathcal{S}.\text{top}()$ if $\mathcal{S}\neq\emptyset$
7:   if $t_{i}=\tau_{\text{guideline}}^{+}$ then $\mathcal{S}.\text{push}(\{p_{\text{end}}:-1,\ell_{\max}:0\})$ $\triangleright$ Open new <guideline> block
8:   else if $t_{i}=\tau_{\text{guideline}}^{-}$ then $b.p_{\text{end}}\leftarrow\mathbf{P}[i]$ $\triangleright$ Mark <guideline> end position
9:   else if $t_{i}=\tau_{\text{step}}^{+}$ and $b.p_{\text{end}}\geq 0$ then $\mathbf{P}[i:]\leftarrow\mathbf{P}[i:]-(\mathbf{P}[i]-b.p_{\text{end}}-1)$ $\triangleright$ Reset to end
10:   else if $t_{i}=\tau_{\text{step}}^{-}$ then $b.\ell_{\max}\leftarrow\max(b.\ell_{\max},\mathbf{P}[i]-b.p_{\text{end}})$ $\triangleright$ Track length of max step
11:   else if $t_{i}=\tau_{\text{parallel}}^{-}$ then $\mathbf{P}[i:]\leftarrow\mathbf{P}[i:]-(\mathbf{P}[i]-b.p_{\text{end}}-b.\ell_{\max}-1)$ $\triangleright$ Align to max
12:     $\mathcal{S}.\text{pop}()$ $\triangleright$ Close <guideline> block
13:  return $\mathbf{P}$
Parallel Attention Mask & Positional Encoding.
To support structured parallel generation, we adopt the core design of Multiverse Attention (multiverse) when constructing both the parallel attention mask and the corresponding positional encoding (Algorithm˜1 and Algorithm˜2). This design enables multiple reasoning paths to coexist within a single forward pass while allowing fast adaptation from only a few examples. It also permits efficient KV-cache reuse for the shared context inside the NPR Engine (§2.5), reducing inference overhead. Furthermore, to ensure the model can emit the required structural tags, we initialize a set of special tokens that correspond to these tags and expose them during the cold-start training stage.
Parallel Warmup.
With the parallel mask and positional scheme in place, we perform a supervised warmup step on the distilled dataset $\mathcal{D}_{\text{accept}}$. The model is trained using standard negative log-likelihood. This stage produces the NPR-Beta, which serve as a stable initialization for the subsequent parallel reinforcement learning stage.
2.4 Stage 3: Native-parallel RL
While Parallel-SFT teaches the model the basic primitives of native parallel reasoning, supervised imitation alone is not sufficient. SFT-distilled trajectories tend to lack structural diversity, and some reasoning modes do not generalize beyond the training distribution. To amplify and generalize these capabilities, we introduce a dedicated Native-Parallel RL stage, as shown in Figure˜3. Because NPR-Beta already learns consistent parallel patterns, it serves as a reliable initialization for direct RL.
Figure 3: Comparison of GRPO-style RL (grpo) and Parallel-Aware Policy Optimization.
Below we summarize the practical modifications we make to standard RL (dapo) to respect parallel semantics and stabilize training.
(1) Parallel rollouts with our parallel inference engine.
Existing inference engines (vllm, sglang) do not enforce strict parallel semantics, so they can produce malformed trajectories. We therefore sample rollouts using our NPR-Engine (§2.5), which guarantees that every generated trajectory follows the intended Map–Process–Reduce flow.
(2) Structural filtering during rollout.
Even with a structured engine, rare format violations can occur. To prevent malformed sequences from entering optimization, we perform schema-level filtering during rollout. Rather than relying solely on a text-based format checker [^2] We found that the text-based format checker always misses rare corner cases., we use the SFT-constructed attention-mask and position-id encoding that exactly represent the parallel schema. After filtering, all retained rollouts strictly obey the target structure; therefore the reward reduces to accuracy only333+1 for a correct final answer, -1 otherwise..
(3) Batch-level advantage normalization.
Because format-violating samples are removed before optimization, group-level variance collapses, which makes relative (group) advantages ineffective. We adopt a Lite-PPO (lite_ppo) style advantage but replace group-level variance with batch-level variance. For each sample $i$ and token $t$ we compute
$\hat{A}_{i,t}:=\frac{R_{i}-\text{mean}\!\left(\{R_{1},R_{2},\cdots,R_{G}\}\right)}{\text{std}\!\left(\{R_{1},R_{2},\cdots,R_{G}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0},\cdots,R_{N*G}}\}\right)},$ (5)
where $N$ is batch size and $G$ is group size, and $R$ is the accuracy reward described above.
(4) Preserve gradients on special tokens.
Special tokens [^4] The tags that control parallel branching and merging. are critical to maintain parallel semantics. Token-level clipping that suppresses gradients for these tokens breaks the learned structure, so we remove clip-masking and ensure special tokens always receive gradients.However, removing clip-masking makes importance-sampling ratios in PPO (ppo) unstable. To avoid unstable reweighting, we eliminate importance sampling and adopt a strict on-policy objective. This both stabilizes training and speeds it up because we do not need to recompute historical log-probabilities.
Putting these choices together yields our Parallel-Aware Policy Optimization (PAPO) objective:
$\mathcal{J}(\theta)=\mathbb{E}_{(q,y)\sim\mathcal{D},\,\{\hat{y}_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot\mid q)}-\frac{1}{\sum_{i=1}^{G}|\hat{y}_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|\hat{y}_{i}|}\big[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{\pi_{\theta}(\hat{y}_{i,t}\mid q,\hat{y}_{i,<t})}{\text{sg}[\pi_{\theta}(\hat{y}_{i,t}\mid q,\hat{y}_{i,<t})]}}\hat{A}_{i,t}\big].$ (6)
where $\text{sg}[\cdot]$ denotes stop-gradient. In practice, the stop-gradient fraction acts to preserve on-policy gradient flow while avoiding unstable importance reweighting.
2.5 Engineering Enhancement: NPR Engine
Multiverse’s parallel-generation engine (multiverse) supplies a powerful substrate for large-scale rollout based on SGLang [^5] https://github.com/sgl-project/sglang, but when exercised at production scale, it exposed a set of brittle implementation corners that undermine both correctness and RL stability. We implemented a compact set of engine-level mitigation to restore deterministic behavior, memory safety, and correct length accounting across high-throughput parallel rollouts, which together form the NPR-Engine used in our Parallel-RL pipeline.
KV-cache double-free and memory corruption.
Under heavy parallel branching, shared radix-tree KV paths were sometimes recycled more than once when the cache exceeded its capacity; the incidence scaled with the branching factor and produced context corruption and, in pathological cases, GPU memory leakage.
Solution We replace opportunistic recycling with an explicit, budget-aware reclamation strategy: when observed KV usage would exceed the preallocated budget we perform an immediate cache flush and deterministic reallocation of the affected blocks.
Underestimated global token budget.
Parallel decoding multiplies aggregate token consumption roughly by the number of branches, but the original accounting tracked only the longest single branch—allowing runs to exceed the configured max_new_tokens.
Solution We extended length accounting to be branch-aware: the engine now records the active branching factor at each expansion and updates a global token ledger accordingly.
Undefined states from illegal parallel schemas.
Certain parallel-branch layouts fell outside the engine’s conditional logic, producing undefined states in rare corner cases.
Solution We add a lightweight pre-branch format validator that enforces a small set of structural invariants before any expansion. These checks are intentionally cheap and conservative, only structurally valid branchings are permitted, so they prevent illegal states with negligible runtime cost.
Local repetition inside <step> blocks.
Fine-grained step streams tended to exhibit local repetition under parallel sampling, which degraded the clarity of stepwise traces.
Solution We apply a mild, selective repetition penalty (coefficient = 1.02) to tokens generated within <step>...</step> contexts while keeping <guideline> and <takeaway> streams penalty-neutral (1.0).
After integrating these fixes into the verl rollout framework, the NPR-Engine exhibited substantially improved determinism, memory stability, and correctness under large-scale parallel RL workloads. Empirical training and evaluation indicate these engine-level remedies are essential: they prevent subtle off-policy artifacts and stabilise optimization when operating at the throughput demanded by production Parallel-RL.
3 Experiments
3.1 Training Setup.
Training Datasets.
We build our experiments on the ORZ dataset (orz), which contains 57k problem–answer pairs. To ensure consistency across all stages of our pipeline, we sample a fixed subset of 8k examples from ORZ and use this for Stage 1 (§2.2), Stage 2 (§2.3), and Stage 3 (§2.4).
Training Details.
Our models are based on Qwen3-4B-Instruct-2507 and Qwen3-4B (non-thinking mode) (DBLP:journals/corr/abs-2505-09388/qwen3). We intentionally avoid the thinking-mode variant because it cannot be trained with standard supervised fine-tuning (§4.6).
We summarize the key configurations for each stage below.
  • Stage 1. We follow the DAPO setup and allow a maximum generation length of 30,000 tokens.
  • Stage 2. Training begins with a learning rate of 1e-6, which is decayed to 5e-7. We apply a weight decay of 0.1.
  • Stage 3. We employ our PAPO together with the NPR engine. The maximum generation length remains 30,000 tokens, and the learning rate is set to 1e-7.
3.2 Evaluation Setup
Evaluation Metrics.
We measure accuracy using avg@k, defined as the expected proportion of correct answers among $k$ generated solutions for each problem. If the model produces $k$ candidate solutions and $c$ of them are correct, the metric reduces to
$\text{avg@}k\;=\;\frac{c}{k},$ (7)
Evaluation Benchmarks.
We assess the effectiveness and generalization ability of NPR across a diverse suite of reasoning benchmarks. For relatively small-scale datasets such as AIME24 (maa_aime2024), AIME25 (maa_aime2025), HMMT25 (balunovic2025matharena), and AMC23 (maa_amc12_problems_solutions), we report avg@8, which better reflects performance when multiple sampled solutions are available. For larger or more heterogeneous benchmarks including OlympiadBench (he-etal-2024-olympiadbench), Minerva-Math (lewkowycz2022solving), ZebraLogic (zebralogic), and MATH500 (math500), we follow the standard single-answer setting and report avg@1.
Compared Baselines.
We compare NPR against a broad set of strong baselines:
  • Open Sequential Reasoners: Qwen2.5-32B-Instruct (yang2024qwen25), Qwen3-4B (DBLP:journals/corr/abs-2505-09388/qwen3) (without thinking mode), and Qwen3-4B-Instruct-2507.
  • Recent Parallel Reasoners: Multiverse (multiverse) models, including Multiverse-32B and our reproduced Multiverse-4B built on Qwen3-4B-Instruct-2507.
  • Sequential Variants: SR-Beta and SR, both trained purely by the sequential reasoning paradigm.
3.3 Overall Reasoning Performance.
[…]
Model Data Train Base A25 A24 H25 OB MvM ZL AMC23 M500 AVG
Q2.5-32B-Inst. - - - 10.4 15.8 3.8 46.4 40.8 43.6 62.8 80.4 38.0
MV-32B s1.1-8k $\rightarrowP S$FT Q2.5-32B-Inst. 45.8 53.8 20.8 48.0 40.0 47.1 72.5 91.8 52.5
Q3-4B-Inst. - - - 47.4 60.0 31.0 64.0 41.2 80.2 92.2 93.4 63.7
MV-4B s1.1-8k $\rightarrowP S$FT Q3-4B-Inst. 42.9 46.7 20.8 38.8 34.9 60.2 75.0 81.6 50.1
NPR-Beta orz-8k Parallel SFT 42.9 50.8 23.3 60.1 41.2 76.1 85.9 91.6 59.0
SR-Beta Sequential SFT 37.1 52.1 22.5 56.3 41.5 72.8 91.6 92.0 58.2
[0.8pt/2pt]
SR orz-8k Sequential RL NPR-Beta 49.2 57.1 26.3 62.2 38.2 78.9 90.9 92.8 62.0
NPR Parallel RL 50.4 63.3 30.8 63.7 43.0 81.7 93.1 93.6 65.0
Q3-4B - - - 19.1 25.0 12.1 48.6 28.5 35.2 65.6 84.8 39.9
NPR-Beta orz-8k Parallel SFT Q3-4B 43.8 52.5 29.2 57.8 45.9 70.0 85.3 86.8 58.9
NPR Parallel RL NPR-Beta 53.8 62.5 32.9 61.9 47.1 75.8 89.7 91.8 64.4
The main experimental results are summarized in Table˜2. Across all benchmarks, NPR demonstrates substantial gains over strong baselines (Qwen3-4B-Instruct-2507 and Qwen3-4B without thinking mode) and consistently outperforms both Multiverse-32B and Multiverse-4B.
Training-data advantage
A key source of improvement comes from replacing the Multiverse training corpus (s1.1-8k for MV-4B) with our self-distilled dataset (orz-8k for NPR-Beta). Although the two pipelines differ slightly in implementation details, both rely on parallel-style SFT, making the comparison meaningful. The impact of the data substitution is clear and consistent: performance on AIME24 increases from 46.7 to 50.8 (+4.1), ZebraLogic from 60.2 to 76.1 (+15.9), AMC23 from 75.0 to 85.9 (+10.9), and MATH500 from 81.6 to 91.6 (+10.0). Overall, the average score improves from 50.1 to 59.0 (+8.9).
Summary These results indicate that our self-distillation corpus produces more accurate and diverse candidate solutions, whereas the Multiverse dataset, which was constructed from sequential reasoning traces, provides limited coverage of genuinely parallel reasoning patterns.
Parallel SFT advantage
Switching from a sequential SFT procedure (e.g., SR-Beta) to our parallel SFT approach (NPR-Beta) leads to consistent improvements across a variety of reasoning benchmarks. Sequential SFT imposes strong step-dependency priors, which limit flexible task decomposition. In contrast, our parallel SFT exposes the model to structurally parallel trajectories during training, enabling more independent subproblem exploration. Concretely, AIME25 improves from 37.1 to 42.9 (+5.8), OlympiadBench from 56.3 to 60.1 (+3.8), HMMT25 from 22.5 to 23.3 (+0.8), and ZebraLogic from 72.8 to 76.1 (+3.3). Overall performance increases from 58.2 to 59.0 (+0.8), with only minor regressions on a few benchmarks.
Summary These findings demonstrate that parallel-format supervision encourages more adaptable and structurally diverse reasoning behaviors, alleviating the restrictive bias inherent in sequential SFT and improving robustness in downstream parallel generation.
Parallel RL advantage
Building on NPR-Beta, applying our parallel RL algorithm yields further gains and consistently surpasses sequential RL (NPR vs. SR). The improvements are broad and systematic: AIME24 rises from 57.1 to 63.3 (+6.2), HMMT25 from 26.3 to 30.8 (+4.5), and Minerva-Math from 38.2 to 43.0 (+4.8). Additional benchmarks show steady gains as well, AIME25 (+1.2), OlympiadBench (+1.5), ZebraLogic (+2.8), AMC23 (+2.2), and MATH500 (+0.8). Overall, the average score increases from 62.0 to 65.0 (+3.0).
Summary The consistent improvements confirm that parallel RL more effectively amplifies high-reward reasoning modes learned during parallel SFT. Our PAPO and stable NPR Engine jointly enable reliable structural exploration and stronger performance across benchmarks.
4 Analyses and Discussion
4.1 Inference Acceleration While Improving Effectiveness
[…]
Table 3: Evaluation results of tokens per second (TPS) and speedup ratio on selected benchmarks. The speedup ratio (denoted as Speedup) is calculated by comparing with sequential reasoning baseline.
Method AIME25 AIME24 HMMT25 AMC23 ZebraLogic
TPS Speedup TPS Speedup TPS Speedup TPS Speedup TPS Speedup
[…]
[0.8pt/2pt]
[…]
4.2 Parallel Reasoning Trigger Analysis
We quantify a model’s tendency to produce simultaneous, non-sequential reasoning using the parallel reasoning trigger rate:
$\text{parallel\_rate}=\frac{N_{\text{parallel}}}{N_{\text{total}}}\times 100\%$ (8)
where $N_{\text{parallel}}$ denotes the number of solutions exhibiting parallel reasoning and $N_{\text{total}}$ the total number of evaluated test cases. Table˜4 reports the parallel rate for the Multiverse baseline (MV-32B) and our NPR model (NPR-Inst.) across eight benchmark sets (AIME25, AIME24, HMMT25, OlympiadBench, Minerva, ZebraLogic, AMC23, and MATH500).
MV-32B displays substantial variability in its parallel rate across datasets, indicating that its adoption of parallel reasoning is highly dataset-dependent. In particular, performance on logic-intensive tasks such as ZebraLogic is markedly lower than on several math contest datasets, suggesting that the Multiverse training paradigm, which gradually transitions from sequential to parallel behavior, yields inconsistent internalization of parallel strategies and is sensitive to domain characteristics.
By contrast, our NPR model attains a uniform 100.0% parallel rate across all eight datasets. This consistency implies that the end-to-end NPR training pipeline more reliably institutionalizes parallel reasoning as the model’s default problem-solving mode, independent of dataset domain or complexity. Practically, this means NPR not only triggers parallel reasoning more often, but does so robustly across heterogeneous evaluation sets.
Table 4: Comparison of parallel reasoning trigger rates between NPR and MultiVerse across datasets.
Model AIME25 AIME24 HMMT25 Olympiad Minerva ZebraLogic AMC23 MATH500
MV-32B 65.0 62.9 63.3 69.5 66.9 45.8 70.0 76.0
NPR-Inst. 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
4.3 Test-time Scalability
We evaluate NPR’s test-time scalability using the avg@8 and best@8 scores reported in Table˜5. The results show that NPR reliably increases oracle coverage at test time, with the largest and most consistent gains occurring when the base model is relatively weak. For the Non-thinking backbone, supervised fine-tuning raises best@8 on AIME25 from 36.7 to 70.0, and NPR further increases it to 76.7, a 6.7 point improvement over SFT. On HMMT25 for the same backbone, best@8 moves from 23.3 to 46.7 after SFT and then to 53.3 with NPR, a further 6.6 points. For the Instruct backbone, NPR raises AIME25 best@8 to 70.0 compared with 63.3 for SFT. Overall, NPR amplifies the coverage benefits introduced by SFT and converts modest increases in sample diversity into meaningful gains in best@8, although the magnitude of improvement depends on the task and the starting strength of the backbone.
Table 5: Performance shown for SFT and RL checkpoints on both the Instruct and Non-thinking Qwen3-4B backbones.
AIME25 AIME24 HMMT25 AMC23
avg@8 best@8 avg@8 best@8 avg@8 best@8 avg@8 best@8
Qwen3-4B-Instruct-2507 47.4 63.3 60.0 86.7 31.0 46.7 92.2 96.7
NPR-Beta-Inst. 42.9 63.3 50.8 83.3 23.3 46.7 85.9 97.5
NPR-Inst. 50.4 70.0 63.3 80.0 30.8 53.3 93.1 100.0
Qwen3-4B-Non-thinking 19.1 36.7 25.0 40.0 12.1 23.3 65.6 93.3
NPR-Beta-Non. 43.8 70.0 52.5 83.3 29.2 46.7 85.3 97.5
NPR-Non. 53.8 76.7 62.5 80.0 32.9 53.3 89.7 100.0
4.4 Evolution Dynamics Toward NPR
As shown in Figure˜4, the evolution toward native parallel reasoning (NPR) is gradual and structured. Naively enforcing the parallel generation format at the outset severely degrades performance (for example, Qwen3-4B-Instruct-2507 falls on AIME25 from 47.5 to 17.5). To address this, we adopt a three-stage pipeline. Stage 1 applies format-following reinforcement learning to stabilize format compliance and correctness, producing reliable trajectories that serve only as training data for the next stage [^6] Stages 1 and 2 are trained from the same initialization; Stage 1 supplies data to Stage 2. Stage 2 performs a parallel warmup via supervised fine-tuning, teaching independent branching and correct special-token usage, this structured learning causes a small, transient performance dip. Finally, Stage 3 uses Native Parallel RL to recover and enhance reasoning quality, yielding final results that surpass the autoregressive baseline. Together, the results show that NPR is not merely a consequence of format supervision but emerges from progressively aligning format, parallel structure, and adaptive policy learning.
Figure 4: Learning dynamics of evaluation on AIME 2025.
4.5 Qualitative Case Studies
Table˜6 and Table˜7 present two representative examples that highlight how our system performs parallel reasoning. Both solutions are generated entirely through the parallel inference pipeline and are answered correctly. A key property of the system is its adaptivity: more challenging problems automatically trigger broader exploration, while easier prompts elicit only a small number of focused branches. The model organizes its computation into parallel blocks, with the final block often functioning as a multi-angle verifier that cross-checks or refines candidate solutions.
Parallelism operates in two complementary ways. First, it can explore multiple independent solution strategies simultaneously. Second, it can decompose a single problem into smaller subproblems that are solved in parallel. For correct examples, at least one branch converges to a valid answer, and the verification block preserves it. Errors, when they occur, typically arise from an incorrect high-level plan that propagates through many branches or from branches that expand into unnecessarily long, error-prone reasoning chains. By exploring diverse trajectories and applying lightweight consistency checks, the parallel architecture provides substantially greater robustness than relying on a single sequential reasoning path.
Table 6: Example of decomposing a problem into multiple subproblems for independent resolution, followed by reflective reasoning in the <takeaway> block.
Case #1
Problem: What is the domain of the function $f(x)=\frac{2-x}{\log(2-\log(x-2))}$, where $\log$ is the base $10$ logarithm function? Express your answer in interval notation.
NPR Generation:
<guideline>
<plan>1: Determine the domain by identifying all values of $x$ for which the denominator is defined and non-zero, and the logarithmic expressions are valid.</plan>
<plan>2: Break down the inner logarithm $\log(x-2)$ and the outer logarithm $\log(2-\log(x-2))$ to find the required domain constraints.</plan>
<plan>3: Ensure that the argument of every logarithm is positive and that the entire denominator is not zero.</plan>
<guideline>
<step>1: $\cdots$, Therefore, $x=12$ must be excluded from the domain.</step>
<step>2: $\cdots$, So the domain is $(2,102)\setminus\{12\}$, or $(2,12)\cup(12,102)$.</step>
<step>3: $\cdots$, Thus, the condition $x\neq 12$ does not introduce any new restriction because $x=12$ is already excluded by $x<10$. So no additional point needs to be removed from the domain.</step>
[…]
<guideline>
<plan>1: Verify the domain by testing values within the proposed intervals.</plan>
<plan>2: Confirm that all conditions (logarithm domains, non-zero denominator) are fully captured.</plan>
<guideline>
<step>1: $\cdots$, All tests confirm that the domain is $(2,12)\cup(12,102)$.</step>
<step>2: $\cdots$, No other points cause issues. The logarithmic functions are defined in the intervals, and no other values make the denominator zero. Thus, the domain is fully described.</step>
<takeaway>All constraints have been verified through both algebraic reasoning and numerical testing. $\cdots$</takeaway>
The domain of the function $f(x)=\frac{2-x}{\log(2-\log(x-2))}$ is $\boxed{(2,12)\cup(12,102)}$.
Table 7: Examples of exploring different methods to independently solve problems and engaging in reflective reasoning within the <step>.
Case #2
Problem: Altitudes $\overline{AD}$ and $\overline{BE}$ of $\triangle ABC$ intersect at $H$. If $\angle BAC=54^{\circ}$ and $\angle ABC=52^{\circ}$, then what is $\angle AHB$?
NPR Generation:
<guideline>
<plan>1: Use the property that in any triangle, the angle between two altitudes intersecting at the orthocenter can be found using the relationship involving the angles of the triangle.</plan>
<plan>2: Apply the fact that in triangle $ABC$, with altitudes $AD$ and $BE$ intersecting at $H$, the quadrilateral $ADHE$ is cyclic, and use angle chasing in this cyclic quadrilateral.</plan>
<plan>3: Use the known result that $\angle AHB=180^{\circ}-\angle C$, where $H$ is the orthocenter of triangle ABCABC.</plan>
<guideline>
<step>1: $\cdots$, Therefore, $\angle AHB=180^{\circ}-74^{\circ}=106^{\circ}$. </step>
<step>2: $\cdots$, but note that $\angle AED=180^{\circ}-\angle ABE=180^{\circ}-90^{\circ}+\angle ABC=90^{\circ}+52^{\circ}=142^{\circ}$? Wait—better approach: in the cyclic quadrilateral formed by the feet and orthocenter, the angle at $H$ can be found via inscribed angle relationships. Alternatively, consider triangle $ABC$, $\cdots$, then $\angle AHB=180^{\circ}-74^{\circ}=106^{\circ}$. </step>
<step>3: $\cdots$, Therefore, $\angle AHB=180^{\circ}-\angle ACB=180^{\circ}-74^{\circ}=106^{\circ}$.</step>
<takeaway>All three approaches lead to the same conclusion: $\cdots$, The reasoning is consistent across different strategies: either using known orthocenter angle properties, cyclic quadrilateral angle chasing, or direct application of the formula. All confirm the same result.</takeaway>
<guideline>
<plan>1: Verify the result by direct angle chasing in quadrilateral $ADHE$. </plan>
<plan>2: Cross-check with triangle angle sum. </plan>
<guideline>
<step>1: $\cdots$, Wait — actually, $\angle ABD$ is not necessarily that. Wait: $\angle ABD$ is part of $\angle ABC=52^{\circ}$. Actually, since $AD$ is an altitude from $A$, $\cdots$, Wait — in triangle $ABC$, $AD\perp BC$, so $D$ lies on $BC$, $BE\perp AC$, so $E$ lies on $AC$. Then quadrilateral $BDHE$ has right angles at $D$ and $E$, so $BDHE$ is cyclic. Therefore, $\angle BHD=180^{\circ}-\angle BEC$? Wait. Actually, in cyclic quadrilateral $BDHE$, $\cdots$, and no contradiction arises from the other approaches, the conclusion stands.</step>
<step>2: $\cdots$, This confirms the earlier conclusion.</step>
<takeaway>All lines of reasoning—whether through known orthocenter properties, cyclic quadrilateral angle chasing, or direct triangle angle sum—lead to the same result: $\angle AHB=106^{\circ}$. The result is consistent, reliable, and internally verified.</takeaway>
The measure of $\angle AHB$ is $\boxed{106^{\circ}}$.
4.6 Preliminary Exploration
Before arriving at our final design, we conducted an extensive series of preliminary explorations to understand how parallel reasoning could be learned in practice. Our goal in presenting these attempts is not to claim that they are fundamentally flawed, but to show that the final method emerged only after multiple iterations and empirical dead ends.
We initially hoped that a thinking-style model could acquire native parallel reasoning directly by learning inside the <think></think> block. However, because such models have already undergone heavy RL training, we observed poor learning dynamics during SFT: special tokens often fragmented (e.g., <conclusion> degraded into <con>), and increasing their loss weight did not help. Despite repeated attempts to stabilize these tokens, the model failed to reliably learn parallel structures inside the thinking block.
Next, we attempted to preserve the original <think></think> content and train only the portion that follows. This required excluding the internal thinking region from the loss, which we implemented by injecting random meaningless tokens inside the block and masking them out. However, this strategy still yielded unstable behavior, suggesting that the internal reasoning space of thinking models is difficult to modify without disrupting RL priors.
These failures led us to abandon the thinking format entirely and move to an instruct model with a simplified output template. We began by unfolding Multiverse’s multi-branch outputs into separate standalone branches and measured their edit distance; the differences were minimal. SFT on either version produced nearly identical results, motivating us to remove unnecessary branching complexity.
During this process, we discovered that some of Multiverse’s “parallel” successes were actually produced through sequential AR generation, because the format included non-parallel outer content and the training schedule transitioned gradually from sequential to parallel. This motivated a cleaner template design that preserves only the special tags essential for genuine parallel reasoning.
Finally, we attempted to replicate the approach on Qwen2.5 series models such as Qwen2.5-7B-Instruct and vanilla Base LLMs. Their poor instruction-following and weak reasoning ability caused the Stage-1 self-distillation to fail badly, especially on AIME-level problems. These observations reinforced our decision to build NPR on top of Qwen3-4B-Instruct and Qwen3-4B Non-Thinking.
5 Related Work
Parallel Reasoning.
Parallel reasoning improves reasoning efficiency and robustness by exploring multiple reasoning paths simultaneously, unlike standard sequential reasoning which is prone to early commitment errors (the “prefix trap”) and lacks self-correction, leading to suboptimal solutions and slow inference due to its strictly step-by-step generation process (DBLP:journals/corr/abs-2510-12164/parallelreasoning). Early methods, such as Best-of-N (DBLP:journals/corr/abs-2110-14168/bon) and Self-Consistency (DBLP:conf/iclr/0002WSLCNCZ23/self-consistency), select the most scored or consistent output from independent paths but are not end-to-end optimized. Search-based approaches like Tree-of-Thought (Tot), Graph-of-Thought (DBLP:conf/aaai/BestaBKGPGGLNNH24/graph-of-thoughts), Monte Carlo Tree Search (DBLP:journals/corr/abs-2405-00451/mcts) further explore reasoning trees but rely on hand-designed structures and external verifiers, limiting flexibility and scalability. To further improve the adaptability and flexibility of parallel reasoning operations, recent work strives through learning approaches. One line of work adopts the SFT paradigm—for example, Multiverse (multiverse), ParaThinker (DBLP:journals/corr/abs-2509-04475/parathinker), and SSFT (DBLP:journals/corr/abs-2510-05132/globeforking)—which guide model learning through parallel reasoning paths derived from the sequential trajectories of more powerful large reasoning models (LRMs). However, such pure imitation limits the model’s ability to discover novel reasoning patterns. Another line of work enhances parallel reasoning capabilities through reinforcement learning (RL), such as APR (DBLP:journals/corr/abs-2504-15466/APR) and Parallel-R1 (parallel_r1). However, these methods either demonstrate effectiveness only on toy tasks or still depend on supervised data from other reasoning models to bootstrap the RL process.
RL for Reasoning.
Reinforcement learning (RL) has become an important tool for enhancing the reasoning capabilities of large language models (LLMs) in recent years (DBLP:journals/corr/abs-2403-04642, DBLP:journals/corr/abs-2509-08827, tokenswift, seek, rulereasoner, routerlens, cream). Early and widely adopted approaches—such as RL from human feedback (RLHF)—optimize outcome-level rewards derived from human preferences or task-level correctness (simpo). These methods improve alignment and robustness in general generation tasks but provide only coarse control over intermediate reasoning trajectories. Then, research shifts toward process-aware RL, where step-level reward modeling offer denser and more interpretable supervision (oai_prm, DBLP:conf/acl/ZhangZWZLYLZL25, DBLP:journals/corr/abs-2504-16828). Those process-level feedback, however, suffer from subjectivity, high annotation cost, and unstable optimization due to ambiguous or unverifiable intermediate signals. A further evolution leads to Reinforcement Learning with Verifiable Reward (RLVR), which replaces opaque reward models with explicit, auditable verifiers (e.g., logical checkers, rule-based graders, or formal validators) (grpo, DBLP:journals/corr/abs-2502-14768, dapo, gspo). Compared with conventional reward modeling, RLVR provides objectivity, reproducibility, and stronger correctness guarantees, making it particularly suited for reasoning tasks where outputs are verifiable (e.g., math, programming, or factual QA). Moreover, RLVR reduces human labeling costs by leveraging deterministic verifiers as reward oracles.
6 Conclusion
This work presents a simple and scalable framework for building a Native Parallel Reasoner that learns adaptive decomposition, diverse parallel planning, and reliable aggregation without relying on external teacher models. By combining self-distilled parallel SFT with agentic parallel RL, our approach produces genuinely parallel reasoning policies rather than simulated or scripted ones. Experiments on eight reasoning benchmarks show consistent improvements over Multiverse datasets, autoregressive training, and direct RL. Our analysis further demonstrates meaningful inference acceleration, stronger test-time scalability, and the absence of pseudo-parallel behavior. Case studies illustrate how the model adapts its parallelism to problem difficulty, enabling structured exploration and robust verification. These results indicate that native parallel reasoning is a promising direction for more general and scalable intelligence.

### v2

@@
-
-marginparsep has been altered.
-topmargin has been altered.
-marginparpush has been altered.
-The page layout violates the ICML style.Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you.
-We’re not able to reliably undo arbitrary changes to the style.
-Please remove the offending package(s), or layout-changing commands and try again.
+Native Parallel Reasoner: Reasoning in Parallelism via Self-Distilled Reinforcement Learning
 Native Parallel Reasoner: Reasoning in Parallelism via
 Self-Distilled Reinforcement Learning
-Tong Wu $^{\dagger}$ Yang Liu*  Jun Bai*  Zixia Jia Shuyi Zhang  Ziyong Lin  Yanting Wang
+[…]
 […]
@@
 1 Introduction
-The advent of super-scale Large Language Models (LLMs), exemplified by Gemini 3 (gemini3), GPT-5 (gpt5), and DeepSeek-V3.2 (deepseek_v3.2), has shifted the frontier of AI from semantic fluency to deep, multi-step agentic reasoning.
-Despite the excitement of “deeper” test-time scaling that enables models to solve complex problems muennighoff2025s1simpletesttimescaling, the “wider” reasoning capacity to explore diverse trajectories in parallel emerges as the dominant requirement toward agentic AI (shen2025thinkingvsdoingagents, gemini2.5).
-The MapReduce paradigm has long underpinned distributed computing by separating task decomposition from trajectory aggregation (mapreduce, multiverse, wang2025llmtimesmapreducev2), yet its application to agentic language modeling remains a critical missing link in the evolution of open-source LLMs.
+The advent of super-scale Large Language Models (LLMs), exemplified by Gemini 3 (Pichai et al., 2025), GPT-5 (OpenAI, 2025), and DeepSeek-V3.2 (DeepSeek, Inc., 2025), has shifted the frontier of AI from semantic fluency to deep, multi-step agentic reasoning.
+Despite the excitement of “deeper” test-time scaling that enables models to solve complex problems Muennighoff et al.
+(2025), the “wider” reasoning capacity to explore diverse trajectories in parallel emerges as the dominant requirement toward agentic AI (Shen et al., 2025, Comanici et al., 2025).
+The MapReduce paradigm has long underpinned distributed computing by separating task decomposition from trajectory aggregation (Dean and Ghemawat, 2008, Yang et al., 2025a, Wang et al., 2025a), yet its application to agentic language modeling remains a critical missing link in the evolution of open-source LLMs.
 Ideally, the model should internalize the collaborative breadth of multi-agent systems directly into an efficient, natively parallel architecture.
@@
 First, Algorithmic and Architectural Incompatibility.
-Prevalent inference engines (sglang, vllm) and Reinforcement Learning (RL) algorithms (grpo) are ill-equipped for native branching: The former fails to control parallel branching and aggregation; the latter often clips the gradients of the special tokens that trigger those operations, preventing the model from learning strict structure.
+Prevalent inference engines (Zheng et al., 2024, Kwon et al., 2023) and Reinforcement Learning (RL) algorithms (Shao et al., 2024) are ill-equipped for native branching: The former fails to control parallel branching and aggregation; the latter often clips the gradients of the special tokens that trigger those operations, preventing the model from learning strict structure.
 Second, Inefficient Hand-Crafted Parallelism.
-Although the intuitive advantage of parallel sampling lies in efficiency, early attempts on internalizing the parallelism parallel_r1, DBLP:journals/corr/abs-2509-04475/parathinker, zhao2025parallelsearch resort to hand-crafted divide-and-conquer rules via independent sampling.
+Although the intuitive advantage of parallel sampling lies in efficiency, early attempts on internalizing the parallelism Zheng et al.
+(2025a), Wen et al.
+(2025), Zhao et al.
+(2025) resort to hand-crafted divide-and-conquer rules via independent sampling.
 These methods fail to leverage shared Key-Value (KV) states, necessitating redundant recalculations for every branch and resulting in prohibitive linear latency costs ($O(N)$) that render the model impractical for real-time deployment.
 Third, Reliance on Supervised Distillation.
-Framework such as Multiverse (multiverse) successfully operationalizes native parallelism but depends heavily on supervised data distilled from stronger teacher models.
+Framework such as Multiverse (Yang et al., 2025a) successfully operationalizes native parallelism but depends heavily on supervised data distilled from stronger teacher models.
 While effective for compressing capabilities into smaller models, this dependence restricts the student to mimicking the teacher’s sequential reasoning topology force-fitted into a parallel format, imposing an "Intelligence Ceiling" that prevents it from novel, model-intrinsic parallel strategies necessary for super-intelligence.
@@
 Specifically, NPR employs a three-stage progressive training paradigm designed to transition the model from sequential emulation to genuine parallel cognition.
-In Stage 1 ($\S$2.2), we warm up a seed instruction-tuned model (e.g., Qwen3-4B-Instruct) to spontaneously discover valid parallel structures by applying standard DAPO (dapo) with a format-aware reward function, yielding a structured trajectory generator, NPR-Zero, which produces parallel-formatted outputs but still relies on sequential visibility, i.e., simulated parallelism.
-In Stage 2 ($\S$2.3), we bridge the gap to native parallel architecture by performing rejection sampling on NPR-Zero and conducting parallel warmup, which instills strict topological constraints using parallel positional encoding and attention mask as in multiverse, converting the sequential behavior of models into real parallel execution.
+In Stage 1 ($\S$2.2), we warm up a seed instruction-tuned model (e.g., Qwen3-4B-Instruct) to spontaneously discover valid parallel structures by applying standard DAPO (Yu et al., 2025) with a format-aware reward function, yielding a structured trajectory generator, NPR-Zero, which produces parallel-formatted outputs but still relies on sequential visibility, i.e., simulated parallelism.
+In Stage 2 ($\S$2.3), we bridge the gap to native parallel architecture by performing rejection sampling on NPR-Zero and conducting parallel warmup, which instills strict topological constraints using parallel positional encoding and attention mask as in Yang et al.
+(2025a), converting the sequential behavior of models into real parallel execution.
 Finally, Stage 3 ($\S$2.4) generalizes these capabilities beyond the initial distribution via Native-Parallel RL.
@@
 The results reveal three consistent superiorities of NPR ($\S$3.3).
-  • Self-Distilled Data Efficacy: Our self-distilled datasets outperform previous teacher-generated trajectories in multiverse by an average of 10.1 points, validating the hypothesis to learn from native distributions.
+  • Self-Distilled Data Efficacy: Our self-distilled datasets outperform previous teacher-generated trajectories in Yang et al.
+(2025a) by an average of 10.1 points, validating the hypothesis to learn from native distributions.
   • Parallelism Effectiveness and Efficiency: Both NPR-Beta and NPR-RL yields significant performance gains over direct sequential RL baselines (e.g., DAPO), confirming that adaptive parallel policies provide a superior search mechanism compared to single-path rollouts.
@@
 […]
-  • We re-engineered an NPR Engine, which redesigns core components of parallel execution, including radix-cache memory management, output-length control, and parallel-state branching based on Multiverse-Engine (multiverse).
+  • We re-engineered an NPR Engine, which redesigns core components of parallel execution, including radix-cache memory management, output-length control, and parallel-state branching based on Multiverse-Engine (Yang et al., 2025a).
 This eliminates the instability inherent in standard engines, providing a practical backend for native parallel RL training.
@@
 Policy Optimization for Language Models
-To optimize the policy model within our reinforcement learning framework, we adopt objective functions based on DAPO (dapo).
+To optimize the policy model within our reinforcement learning framework, we adopt objective functions based on DAPO (Yu et al., 2025).
 We first introduce the original DAPO update procedure.
@@
 2.2 Stage 1: Format-follow Reinforcement Learning
-To support adaptive decomposition and parallel reasoning during generation, we adopt a simplified “Map–Process–Reduce” schema inspired by Multiverse (multiverse) but with a leaner structure.
+To support adaptive decomposition and parallel reasoning during generation, we adopt a simplified “Map–Process–Reduce” schema inspired by Multiverse (Yang et al., 2025a) but with a leaner structure.
 Each parallel block begins with <guideline> ...
@@
 While this schema provides a clear, learnable format for parallel reasoning, obtaining large-scale, high-quality training data for it remains challenging.
-Prior work such as Multiverse (multiverse) constructs large, multi-step synthetic pipelines and aggregates outputs from several state-of-the-art teacher models (e.g., Deepseek R1 (deepseek_r1) and Gemini 2.5 Pro (gemini2.0)) to overcome data scarcity.
+Prior work such as Multiverse (Yang et al., 2025a) constructs large, multi-step synthetic pipelines and aggregates outputs from several state-of-the-art teacher models (e.g., Deepseek R1 (Guo et al., 2025) and Gemini 2.5 Pro (Google, 2025)) to overcome data scarcity.
 While effective, these multi-teacher pipelines add operational complexity, require access to strong external teachers, and incur substantial maintenance costs.
 We adopt a simpler, self-improving approach.
-Starting from a single pretrained LLM, we apply DAPO (dapo) to induce the target native parallel-reasoning generation format without paired supervision or external teachers.
+Starting from a single pretrained LLM, we apply DAPO (Yu et al., 2025) to induce the target native parallel-reasoning generation format without paired supervision or external teachers.
 Our reward function combines format and accuracy signals.
@@
 Applying this filter yields the distilled dataset
-$\begin{split}\mathcal{D}_{\text{accept}}=\{(q_{i},r^{i}_{j},a^{i}_{j})\mid i\leq N,j\leq K,\text{ s.t.
-}(r^{i}_{j},a^{i}_{j})\sim\pi_{\theta}(\cdot|q_{i}),\mathbb{1}_{\text{accept}}(r^{i}_{j},\hat{a}^{i}_{j})=1\}.\end{split}$ (4)
+$\begin{split}\mathcal{D}_{\text{accept}}=\{(q_{i},r^{i}_{j},\hat{a}^{i}_{j})\mid i\leq N,j\leq K,\text{ s.t.
+}(r^{i}_{j},\hat{a}^{i}_{j})\sim\pi_{\theta}(\cdot|q_{i}),\mathbb{1}_{\text{accept}}(r^{i}_{j},\hat{a}^{i}_{j})=1\}.\end{split}$ (4)
 These accepted trajectories serve as the training corpus for the subsequent supervised fine-tuning stage, which provides a stable initialization for the parallel RL procedure described in §2.4.
@@
 Parallel Attention Mask & Positional Encoding.
-To support structured parallel generation, we adopt the core design of Multiverse Attention (multiverse) when constructing both the parallel attention mask and the corresponding positional encoding (Algorithm˜1 and Algorithm˜2).
+To support structured parallel generation, we adopt the core design of Multiverse Attention (Yang et al., 2025a) when constructing both the parallel attention mask and the corresponding positional encoding (Algorithm˜1 and Algorithm˜2).
 This design enables multiple reasoning paths to coexist within a single forward pass while allowing fast adaptation from only a few examples.
@@
 Because NPR-Beta already learns consistent parallel patterns, it serves as a reliable initialization for direct RL.
-Figure 3: Comparison of GRPO-style RL (grpo) and Parallel-Aware Policy Optimization.
-Below we summarize the practical modifications we make to standard RL (dapo) to respect parallel semantics and stabilize training.
+Figure 3: Comparison of GRPO-style RL (Shao et al., 2024) and Parallel-Aware Policy Optimization.
+Below we summarize the practical modifications we make to standard RL (Yu et al., 2025) to respect parallel semantics and stabilize training.
 (1) Parallel rollouts with our parallel inference engine.
-Existing inference engines (vllm, sglang) do not enforce strict parallel semantics, so they can produce malformed trajectories.
+Existing inference engines (Kwon et al., 2023, Zheng et al., 2024) do not enforce strict parallel semantics, so they can produce malformed trajectories.
 We therefore sample rollouts using our NPR-Engine (§2.5), which guarantees that every generated trajectory follows the intended Map–Process–Reduce flow.
@@
 Because format-violating samples are removed before optimization, group-level variance collapses, which makes relative (group) advantages ineffective.
-We adopt a Lite-PPO (lite_ppo) style advantage but replace group-level variance with batch-level variance.
+We adopt a Lite-PPO (Liu et al., 2025a) style advantage but replace group-level variance with batch-level variance.
 For each sample $i$ and token $t$ we compute
@@
 are critical to maintain parallel semantics.
-Token-level clipping that suppresses gradients for these tokens breaks the learned structure, so we remove clip-masking and ensure special tokens always receive gradients.However, removing clip-masking makes importance-sampling ratios in PPO (ppo) unstable.
+Token-level clipping that suppresses gradients for these tokens breaks the learned structure, so we remove clip-masking and ensure special tokens always receive gradients.However, removing clip-masking makes importance-sampling ratios in PPO (Schulman et al., 2017) unstable.
 To avoid unstable reweighting, we eliminate importance sampling and adopt a strict on-policy objective.
@@
 2.5 Engineering Enhancement: NPR Engine
-Multiverse’s parallel-generation engine (multiverse) supplies a powerful substrate for large-scale rollout based on SGLang [^5] https://github.com/sgl-project/sglang, but when exercised at production scale, it exposed a set of brittle implementation corners that undermine both correctness and RL stability.
+Multiverse’s parallel-generation engine (Yang et al., 2025a) supplies a powerful substrate for large-scale rollout based on SGLang [^5] https://github.com/sgl-project/sglang, but when exercised at production scale, it exposed a set of brittle implementation corners that undermine both correctness and RL stability.
 We implemented a compact set of engine-level mitigation to restore deterministic behavior, memory safety, and correct length accounting across high-throughput parallel rollouts, which together form the NPR-Engine used in our Parallel-RL pipeline.
@@
 Training Datasets.
-We build our experiments on the ORZ dataset (orz), which contains 57k problem–answer pairs.
+We build our experiments on the ORZ dataset (Hu et al., 2025), which contains 57k problem–answer pairs.
 To ensure consistency across all stages of our pipeline, we sample a fixed subset of 8k examples from ORZ and use this for Stage 1 (§2.2), Stage 2 (§2.3), and Stage 3 (§2.4).
 Training Details.
-Our models are based on Qwen3-4B-Instruct-2507 and Qwen3-4B (non-thinking mode) (DBLP:journals/corr/abs-2505-09388/qwen3).
+Our models are based on Qwen3-4B-Instruct-2507 and Qwen3-4B (non-thinking mode) (Yang et al., 2025b).
 We intentionally avoid the thinking-mode variant because it cannot be trained with standard supervised fine-tuning (§4.6).
@@
 We assess the effectiveness and generalization ability of NPR across a diverse suite of reasoning benchmarks.
-For relatively small-scale datasets such as AIME24 (maa_aime2024), AIME25 (maa_aime2025), HMMT25 (balunovic2025matharena), and AMC23 (maa_amc12_problems_solutions), we report avg@8, which better reflects performance when multiple sampled solutions are available.
-For larger or more heterogeneous benchmarks including OlympiadBench (he-etal-2024-olympiadbench), Minerva-Math (lewkowycz2022solving), ZebraLogic (zebralogic), and MATH500 (math500), we follow the standard single-answer setting and report avg@1.
+For relatively small-scale datasets such as AIME24 (Mathematical Association of America, 2024), AIME25 (Mathematical Association of America, 2025), HMMT25 (Balunović et al., 2025), and AMC23 (Mathematical Association of America, 2023), we report avg@8, which better reflects performance when multiple sampled solutions are available.
+For larger or more heterogeneous benchmarks including OlympiadBench (He et al., 2024), Minerva-Math (Lewkowycz et al., 2022), ZebraLogic (Lin et al., 2025), and MATH500 (Hendrycks et al., 2021), we follow the standard single-answer setting and report avg@1.
 Compared Baselines.
 We compare NPR against a broad set of strong baselines:
-  • Open Sequential Reasoners: Qwen2.5-32B-Instruct (yang2024qwen25), Qwen3-4B (DBLP:journals/corr/abs-2505-09388/qwen3) (without thinking mode), and Qwen3-4B-Instruct-2507.
-  • Recent Parallel Reasoners: Multiverse (multiverse) models, including Multiverse-32B and our reproduced Multiverse-4B built on Qwen3-4B-Instruct-2507.
+  • Open Sequential Reasoners: Qwen2.5-32B-Instruct (Qwen et al., 2024), Qwen3-4B (Yang et al., 2025b) (without thinking mode), and Qwen3-4B-Instruct-2507.
+  • Recent Parallel Reasoners: Multiverse (Yang et al., 2025a) models, including Multiverse-32B and our reproduced Multiverse-4B built on Qwen3-4B-Instruct-2507.
   • Sequential Variants: SR-Beta and SR, both trained purely by the sequential reasoning paradigm.
@@
 Parallel Reasoning.
-Parallel reasoning improves reasoning efficiency and robustness by exploring multiple reasoning paths simultaneously, unlike standard sequential reasoning which is prone to early commitment errors (the “prefix trap”) and lacks self-correction, leading to suboptimal solutions and slow inference due to its strictly step-by-step generation process (DBLP:journals/corr/abs-2510-12164/parallelreasoning).
-Early methods, such as Best-of-N (DBLP:journals/corr/abs-2110-14168/bon) and Self-Consistency (DBLP:conf/iclr/0002WSLCNCZ23/self-consistency), select the most scored or consistent output from independent paths but are not end-to-end optimized.
-Search-based approaches like Tree-of-Thought (Tot), Graph-of-Thought (DBLP:conf/aaai/BestaBKGPGGLNNH24/graph-of-thoughts), Monte Carlo Tree Search (DBLP:journals/corr/abs-2405-00451/mcts) further explore reasoning trees but rely on hand-designed structures and external verifiers, limiting flexibility and scalability.
+Parallel reasoning improves reasoning efficiency and robustness by exploring multiple reasoning paths simultaneously, unlike standard sequential reasoning which is prone to early commitment errors (the “prefix trap”) and lacks self-correction, leading to suboptimal solutions and slow inference due to its strictly step-by-step generation process (Wang et al., 2025b).
+Early methods, such as Best-of-N (Cobbe et al., 2021) and Self-Consistency (Wang et al., 2023), select the most scored or consistent output from independent paths but are not end-to-end optimized.
+Search-based approaches like Tree-of-Thought (Yao et al., 2023), Graph-of-Thought (Besta et al., 2024), Monte Carlo Tree Search (Xie et al., 2024) further explore reasoning trees but rely on hand-designed structures and external verifiers, limiting flexibility and scalability.
 To further improve the adaptability and flexibility of parallel reasoning operations, recent work strives through learning approaches.
-One line of work adopts the SFT paradigm—for example, Multiverse (multiverse), ParaThinker (DBLP:journals/corr/abs-2509-04475/parathinker), and SSFT (DBLP:journals/corr/abs-2510-05132/globeforking)—which guide model learning through parallel reasoning paths derived from the sequential trajectories of more powerful large reasoning models (LRMs).
+One line of work adopts the SFT paradigm—for example, Multiverse (Yang et al., 2025a), ParaThinker (Wen et al., 2025), and SSFT (Jia et al., 2025)—which guide model learning through parallel reasoning paths derived from the sequential trajectories of more powerful large reasoning models (LRMs).
 However, such pure imitation limits the model’s ability to discover novel reasoning patterns.
-Another line of work enhances parallel reasoning capabilities through reinforcement learning (RL), such as APR (DBLP:journals/corr/abs-2504-15466/APR) and Parallel-R1 (parallel_r1).
+Another line of work enhances parallel reasoning capabilities through reinforcement learning (RL), such as APR (Pan et al., 2025) and Parallel-R1 (Zheng et al., 2025a).
 However, these methods either demonstrate effectiveness only on toy tasks or still depend on supervised data from other reasoning models to bootstrap the RL process.
 RL for Reasoning.
-Reinforcement learning (RL) has become an important tool for enhancing the reasoning capabilities of large language models (LLMs) in recent years (DBLP:journals/corr/abs-2403-04642, DBLP:journals/corr/abs-2509-08827, tokenswift, seek, rulereasoner, routerlens, cream).
-Early and widely adopted approaches—such as RL from human feedback (RLHF)—optimize outcome-level rewards derived from human preferences or task-level correctness (simpo).
+Reinforcement learning (RL) has become an important tool for enhancing the reasoning capabilities of large language models (LLMs) in recent years (Havrilla et al., 2024, Zhang et al., 2025a, Wu et al., 2025, Li et al., 2025, Liu et al., 2025b, Bai et al., 2025, Wu et al., 2024).
+Early and widely adopted approaches—such as RL from human feedback (RLHF)—optimize outcome-level rewards derived from human preferences or task-level correctness (Meng et al., 2024).
 These methods improve alignment and robustness in general generation tasks but provide only coarse control over intermediate reasoning trajectories.
-Then, research shifts toward process-aware RL, where step-level reward modeling offer denser and more interpretable supervision (oai_prm, DBLP:conf/acl/ZhangZWZLYLZL25, DBLP:journals/corr/abs-2504-16828).
+Then, research shifts toward process-aware RL, where step-level reward modeling offer denser and more interpretable supervision (Lightman et al., 2024, Zhang et al., 2025b, Khalifa et al., 2025).
 Those process-level feedback, however, suffer from subjectivity, high annotation cost, and unstable optimization due to ambiguous or unverifiable intermediate signals.
-A further evolution leads to Reinforcement Learning with Verifiable Reward (RLVR), which replaces opaque reward models with explicit, auditable verifiers (e.g., logical checkers, rule-based graders, or formal validators) (grpo, DBLP:journals/corr/abs-2502-14768, dapo, gspo).
+A further evolution leads to Reinforcement Learning with Verifiable Reward (RLVR), which replaces opaque reward models with explicit, auditable verifiers (e.g., logical checkers, rule-based graders, or formal validators) (Shao et al., 2024, Xie et al., 2025, Yu et al., 2025, Zheng et al., 2025b).
 Compared with conventional reward modeling, RLVR provides objectivity, reproducibility, and stronger correctness guarantees, making it particularly suited for reasoning tasks where outputs are verifiable (e.g., math, programming, or factual QA).